# Developing a Thailand solar irradiance map using Himawari-8 satellite imageries and deep learning models

### v1

@@ Abstract
This paper presents an online platform showing Thailand solar irradiance map every 30 minutes available at https://www.cusolarforecast.com. The methodology for estimating global horizontal irradiance (GHI) across Thailand relies on cloud index extracted from Himawari-8 satellite imagery, Ineichen clear-sky model with locally-tuned Linke turbidity, and machine learning models. The methods take clear-sky irradiance, cloud index, re-analyzed GHI and temperature data from the MERRA-2 database, and date-time as inputs for GHI estimation models, including LightGBM, LSTM, Informer, and Transformer. These are benchmarked with the estimate from the SolCast service by evaluation of 15-minute ground GHI data from 53 ground stations over 1.5 years during 2022-2023. The results show that the four models have competitive performances and outperform the SolCast service. The best model is LightGBM with an MAE of 78.58 and RMSE of 118.97 . Obtaining re-analyzed MERRA-2 data for the whole Thailand region is not economically feasible for deployment. When removing these features, the Informer model has a winning performance in MAE of 78.67 . The obtained performance aligns with existing literature by taking the climate zone and time granularity of data into consideration. As the map shows an estimate of GHI over 93,000 grids with a frequent update, the paper also describes a computational framework for displaying the entire map. It tests the runtime performance of deep learning models in the GHI estimation process.
@@ 1 Introduction
SolarAnywhere provides a core technique called the SUNY model developed in the works from [PH13, PIM+02, PSH+15, PSK+18] that uses satellite imagery and numerical weather predictions to better calculate irradiance under cloudy conditions. SolarGIS by [SC10] uses a semi-empirical solar radiation model with data from satellites for explaining cloud properties. The service also provides site adaptation of Solargis data to improve the quality using ground measurements. PVWatts by NREL;see [Dob13], has an interactive page where the user can input a desired location and PV panel’s parameters to calculate monthly solar irradiance and PV power. [So24] offers solar radiation maps in 15-min resolution. The method relies on HelioClim-3 persistence forecast of a clear-sky index and the use of the Meteosat satellite. SolCast service, detailed in [Bri19], provides solar irradiance forecasts and historical data in several time resolutions. It uses GOES-17, which covers the continental US, and Himawari-8 satellite, which covers Asia and the Pacific, to gather information about cloud opacity and weather conditions. The methodology consists of the REST2v5 clear-sky model, the clear-sky detection model (the BRIGHT-SUN), satellite image processing (cloud repositioning, removing artifacts), and computing a polynomial relationship between and cloud pixel brightness. The validation of Solcast was presented in all climate conditions (equatorial, arid, temperate, and cold) and sky conditions (clear, cloudy, overall), where the equatorial condition had the widest spread of GHI estimates (RMSE = 115 in clouded-sky condition) due to towering cumulonimbus cloud during storms that make it hardest to find the relationship of and cloud pixel brightness. The service also estimates PV power when users provide PV metadata such as location, capacity, and panel tilt angle.
@@
While these aforementioned companies provide a wide range of services and use advanced algorithms with several sources of input data, it boils down to the cost of acquiring these solar data for a regional application as the cost often varies with the number of sites. As of March 2024, Thailand has 552 on-grid solar power plants with total capacity of 2,640 MW, shown in Figure 1 (a). Among these solar sites, Figure 1 (b) shows 53 solar sites fully equipped with groud measurements that participates in our study. To accommodate solar energy applications in Thailand in an operational setting, we aim to develop our own solar map that frequently estimates satellite-derived GHI across the country using satellite imagery and local ground measurement. The cost of acquiring irradiance estimates across Thailand (93,061 grids) through commercial services in an operational setting (every 30-minute) is high as compared to a cost of developing our own platform. For this reason, we study literature of satellite-derived GHI estimation techniques in the following sections.
@@ 1.1 Physic-based and empirical methods
Methods in this group quantify how much clouds and other particles in the atmosphere are affecting the incoming solar radiation by compensating with a clear-sky scenario. This relationship is often parametrized in physical parameters such as aerosol optical depth (AOD), reflectance, surface albedo, which can be derived from satellite imagery. The Heliosat model is a semi-empirical method with a core principle that combines a clear-sky model with a cloud index through a piecewise-polynomial relationship. [Dag05] used Heliosat algorithm and calculated a new cloud index from reflectivity derived from the Meteosat satellite where a modification was made on computing the ground reflectivity to save computing time. [KBSH18] extracted the brightness temperature difference from various bands of INSAT-3-D satellite to calculate all-day index as information of cloud, dust, and fog. The GHI estimation relied on the Heliosat-2 model and validated on ground measurements in India. The RMSE ranged from 48 to 128 under sky conditions with dust and fog. [VPN+23] presented an improvement over the MAGIC-Heliosat method on calculating apparent reflectance to compute the cloud index. The results were validated with ground stations from the baseline surface radiation network (BSRN) in Australia and Japan regions. The method’s performance varied according to evaluation on hourly, daily, and monthly averaged data with RMSE of 161, 80, and 70 , respectively.
While the Heliosat model is a well-known method, there are several other semi-empirical approaches for GHI estimation. [QWL+18] showed that Yang’s hybrid model (YHM) performed best with RMSE of 27.94 when comparing to ANN on datasets in China. The physical method YHM relies on the radiation-damping processes, e.g., Rayleigh scattering, aerosol extinction, absorption by ozone, water vaport and permanent gas. These atmospheric variables were derived from MODIS satellite. [QHM+19] presented a mathematical formula of GHI as a function of irradiance with zero surface albedo, global reflectance function of atmosphere and surface albebo, which can be derived from Himawari-8 satellite. The results were validated on 3-year ground station data in Australia with MAE of 55.8 and RMSE of 91.7 . The GHI model with explicit surface albedo was later used in [QHM+21] but all irradiance components were calculated by pre-computing Look Up Table and using radiative transfer model (RTM). The GHI estimation results were validated on data in Australia with an MAE of 54.7 and an RMSE of 90.3 . [BS21] computed the cloud indices using the surface albedo extracted from eight-band imagery of Himawari-8. These cloud indices were converted to clear-sky indices (for all eight bands), and the eight-band clear-sky indices were optimally weighted. The results were validated using data in Singapore, where the scatter plot between the ground-observed clear-sky index and the satellite-estimated values was improved.
@@ 1.2 Site-adaptation techniques
Neural network techniques, which will be presented in the next section, can capture a complex relationship between the satellite data and the corresponding ground-based GHI measurements.
@@ 1.3 Neural network techniques
A methodology that exploits deep-learning or tree-based models to capture a complex relationship between satellite cloud and GHI. To find the best model, we benchmark LSTM, Informer, Transformer, and LightGBM (LGBM) for GHI estimation. Additionally, we improved Informer and Transformer particularly for this application. For LSTM and LGBM, we study the impact of input features. The results show that all these models provide more accurate GHI estimates than the one from a commercial service, SolCast. Our implementation are provided at https://github.com/energyCUEE/solarmap-DL-2024.
@@ 2 Thailand solar map
Our Thailand solar map platform is currently running at https://cusolarforecast.com/, where the methodology is a machine learning model that takes input of cloud index, geographical location, and clear-sky irradiance to produce the GHI estimate. The random forest model had been implemented during 2022 to June 2024. Our continuing plan of updating the model has been discussed in [SWH+24], and the goal of this paper is to improve the model, considering other tree-based and deep-learning models. The service is currently running under the deep learning model, the Informer, as to be presented subsequently in Section 4.3. This section describes the map requirement and the available input data resources for developing the GHI estimation model.
@@ 3.2 Satellite-derived GHI estimation
When training a model, it is natural to enforce the similarity between the prediction and the ground truth . In this study, we benchmark the following methods: LightGBM (LGBM), LSTM, Informer, Transformer, and SolCast service. For SolCast predictions, we have downloaded GHI estimates at 53 stations during the time period of this study. We modified Transformer and Informer to acquire time-series data for solar irradiance forecasting, while building upon existing LSTMs with additional linear layers. The hyper-parameters of these models were tuned and set with the values shown in Table 1.
@@ 4.1 Data preparation
The extracted cloud index and ground GHI data are merged into a synchronized 15-min resolution during 03:00-17:00 (local time in Thailand), totaling 1,195,257 samples or 20,924 days. The target forecasting hours ranges from 7:00 to 17:00. The data is divided with the ratio of 60:20:20 for train:validation:test, where each of the three sets are ensured to contain a uniform mix of three sky conditions see Table 2. To achieve this, we first segmented the data with lags and stacked them as sequences labeled with sky condition. Then, we shuffle the data and split them for training, validation, and testing.
@@ 4.3 Satellite-derived GHI estimation
Figure 11 shows a comparative performance of all GHI estimation models, including LGBM, LSTM, Informer, Transformer, and the SolCast service. All the proposed benchmarked models provide a similar MAEs, that are 20, 78, 98 in clear, partly cloudy, and cloudy sky conditions, respectively, except for the MAE of SolCast that is the highest in all conditions; see Figure 11 (Top.). For the MBE, LGBM and SolCast have a relatively high negative bias (underestimate) in clear and partly cloudy and with positive bias (overestimate) in the cloudy condition. Meanwhile, LSTM, Informer, and Transformer tend to overestimate—low positive bias in clear and partly cloudy and high positive bias in the cloudy condition, as shown in Figure 11 (Bottom.) The evaluation with the RMSE metric is reported in Table 3. All the proposed models perform almost equally if MAE is considered. However, considering RMSE performance, LGBM has the lowest error in all sky-condition settings. This is because the training loss of the tree model is MSE, while the training loss of deep learning models is set to MAE loss.
| Model | LGBM | LSTM | Informer | Transformer | SolCast |
@@
Our GHI estimation at a half-hourly resolution provides the RMSE values (31.54 - 145.74 ) that aligns with previous studies in equatorial regions, i.e., [VPN+23] reported reported hourly GHI RMSEs of 115-180 across six locations (SAP, TAT, FUA, ISH, DWN, and ASP) using Himawari-8 data; and, [QHM+21] reported an instantaneous GHI RMSE of 118.3 at the DWN station, also using Himawari-8. Our results are in a similarly trend to [KBSH18], providing half-hourly GHI RMSEs of 41.7-167 in South Asia (India). [Bri19] reported an average hourly RMSE of 99.5 W/m² in equatorial regions for zenith angles between 0 and 85 degrees (approx. 05:56-18:37), exceeding our target forecasting hours (7:00-17:00). Note that including GHI estimates from early morning and late evening (nearly night-time) in the analysis could potentially underestimate the MAE and RMSE.
@@
The solar irradiance estimation by Informer is provided in Figure 12. The discrepancies between the estimates and ground truth appear to be larger when the sky is cloudy, as shown in Figure 12 (a). Meanwhile, the MAE evaluated for each hour of the day indicates that the model usually suffers during the daytime (10:00-15:00) under the presence of cloud, as shown in Figure 12 (b). The MAE could go up to 140 in this period, and the positive MBE explains that the model mostly overestimates in this period. Figure 12 (c) and (d) show the selected time series with the top three lowest and highest daily average MAE, respectively. The time series with the lowest error often co-occurs with the clear sky condition where the cloud index is quite low. The time series with the highest error typically happens with cloud fluctuation during the day. Thus, the model functioned as expected; if CI increases, the estimate is decreased. A high estimation error, for example, the middle of plot (d), could come from approximation in the data processing step. It could be the case that CI was missing before 14:00, and the displayed values were interpolated (as a constant of 0.3 from backward imputation).
@@ 4.4 Solar energy in Thailand
Figure 13(c) illustrates hourly-updated GHI estimates on June 15, 2023. Our GHI estimation model can provide solar energy estimation at the finer temporal resolution, compared to [JMPL13] that was aimed for an analysis of a monthly resolution. Our model prioritizes near real-time updates which offers a significant advantage for operational setting in energy applications where real-time update is crucial.
@@ 4.5 Deployment and runtime
Our GHI models are deployed to provide the solar map at two spatial resolutions: zoom level 12 (6,843 grids, approximately per grid) and zoom level 14 (93,061 grids, approximately per grid). To efficiently estimate solar irradiance for all grid points simultaneously, we set the batch size equal to the total number of grids in each map. The runtime performance of LSTM, Informer, and Transformer models on an RTX4060 GPU for this process is presented in Table 5. From Section 3.2.1, when directly including latitude and longitude as features for LGBM, the generated solar map displays some abrupt changes in estimated GHI at some region boundary. These unrealistic results can be fixed when removing these two features, with the cost of having 2-3 MAE increased. For this reason, LGBM results are excluded from the runtime report. All models show acceptable performance ( secs. for the largest batch size), with LSTM being the fastest. Transformer can be an alternative to Informer for their similar performance.
@@ 5 Conclusions
This research presented a methodology of a satellite-derived GHI estimation across Thailand with a spatial resolution of , totaling 93,061 grids with 30-minute updates using cloud indices extracted from Himawari-8 satellite imageries, along with the output of Ineichen and Perez’s clear-sky irradiance model as main features of our GHI estimation models. The study explored recently effective machine learning models, including LightGBM (LGBM), LSTM, Informer, and Transformer, to capture complex cloud effects on ground irradiance attenuation. To optimize computational efficiency, input data is processed as time series, with necessary modifications to Informer and Transformer. Benchmarking against SolCast data from 53 ground stations demonstrates superior performance, with lower MAE and RMSE values. Our GHI estimates, at a half-hourly resolution and with RMSE values ranging from 31.54 to 142.42 W/m², align with previous studies [KBSH18, Bri19, QHM+21, VPN+23]. We provide real-time solar energy maps at hourly, monthly, and yearly resolutions, and assess computational efficiency at different spatial scales. Our monthly results on March, April, and May are aligned with the monthly average of daily global irradiance in year 2013 during the same months [JMPL13]. Additionally, we confirmed the feasibility of using these models for real-time GHI estimation at the required temporal resolution.

### v2

@@ Abstract
This paper presents an online platform showing Thailand solar irradiance map every 30 minutes available at https://www.cusolarforecast.com. The methodology for estimating global horizontal irradiance (GHI) across Thailand relies on cloud index extracted from Himawari-8 satellite imagery, Ineichen clear-sky model with locally-tuned Linke turbidity, and machine learning models. The methods take clear-sky irradiance, cloud index, re-analyzed GHI and temperature data from the MERRA-2 database, and date-time as inputs for GHI estimation models, including LightGBM, LSTM, Informer, and Transformer. These are benchmarked with the estimate from a commercial service X by evaluation of 15-minute ground GHI data from 53 ground stations over 1.5 years during 2022-2023. The results show that the four models have competitive performances and outperform the service X. The best model is LightGBM with an MAE of 78.58 and RMSE of 118.97 . Obtaining re-analyzed MERRA-2 data for the whole Thailand region is not economically feasible for deployment. When removing these features, the Informer model has a winning performance in MAE of 78.67 . The obtained performance aligns with existing literature by taking the climate zone and time granularity of data into consideration. As the map shows an estimate of GHI over 93,000 grids with a frequent update, the paper also describes a computational framework for displaying the entire map. It tests the runtime performance of deep learning models in the GHI estimation process.
@@ 1 Introduction
SolarAnywhere provides a core technique called the SUNY model developed in the works from [PH13, PIM+02, PSH+15, PSK+18] that uses satellite imagery and numerical weather predictions to better calculate irradiance under cloudy conditions. SolarGIS by [SC10] uses a semi-empirical solar radiation model with data from satellites for explaining cloud properties. The service also provides site adaptation of SolarGIS data to improve the quality using ground measurements. PVWatts by NREL;see [Dob13], has an interactive page where the user can input a desired location and PV panel’s parameters to calculate monthly solar irradiance and PV power. [So24] offers solar radiation maps in 15-min resolution. The method relies on HelioClim-3 persistence forecast of a clear-sky index and the use of the Meteosat satellite. SolCast service, detailed in https://www.solcast.com/irradiance-data-methodology, provides solar irradiance forecasts and historical data in several time resolutions. It uses NOAA, EUMETSAT which covers the continental US, Europe, and Himawari-8 satellite from JMA, which covers Asia and the Pacific, to gather information about cloud opacity and weather conditions. The methodology of producing GHI estimates consists of the SolCast cloud model and the REST2v5 clear-sky model. The SolCast cloud model takes historical and forecast data of geostationary meteorological satellite (GMS) and NWP inputs for 14 days ahead. The cloud opacity extracted from satellite cloud detection algorithms is combined with NWP inputs to account for mixing effects from weather artifacts. The validation of Solcast was presented in https://www.solcast.com/validation-and-accuracy for all climate conditions where sites are grouped by continent. The overall normalized MAE and RMSE of SolCast 2023 based on all 207 sites were and , respectively. The service also provides the SolCast Rooftop PV model, which estimates PV power when users provide PV metadata such as location, capacity, and panel tilt angle.
@@
While these aforementioned companies provide a wide range of services and use advanced algorithms with several sources of input data, it boils down to the cost of acquiring these solar data for a regional application as the cost often varies with the number of sites. As of March 2024, Thailand has 552 on-grid solar power plants with a total capacity of 2,640 MW, shown in Figure 1 (a). Among these solar sites, Figure 1 (b) shows 53 solar sites fully equipped with ground measurements that participate in our study. To accommodate solar energy applications in Thailand in an operational setting, we aim to develop our solar map that frequently estimates satellite-derived GHI across the country using satellite imagery and local ground measurements. The cost of acquiring irradiance estimates across Thailand (93,061 grids) through commercial services in an operational setting (every 30 minutes) is high as compared to the cost of developing our own platform. For this reason, we study the literature on satellite-derived GHI estimation techniques in the following sections.
@@ 1.1 Physic-based and empirical methods
Methods in this group quantify how much clouds and other particles in the atmosphere affect the incoming solar radiation by compensating with a clear-sky scenario. This relationship is often parametrized in physical parameters such as aerosol optical depth (AOD), reflectance, surface albedo, which can be derived from satellite imagery. The Heliosat model is a semi-empirical method with a core principle that combines a clear-sky model with a cloud index through a piecewise-polynomial relationship. [Dag05] used Heliosat algorithm and calculated a new cloud index from reflectivity derived from the Meteosat satellite where a modification was made on computing the ground reflectivity to save computing time. [KBSH18] extracted the brightness temperature difference from various bands of INSAT-3-D satellite to calculate the all-day index as information of cloud, dust, and fog. The GHI estimation relied on the Heliosat-2 model and was validated on ground measurements in India. The RMSE ranged from 48 to 128 under sky conditions with dust and fog. [VPN+23] presented an improvement over the MAGIC-Heliosat method on calculating apparent reflectance to compute the cloud index. The results were validated with ground stations from the baseline surface radiation network (BSRN) in Australia and Japan. The method’s performance varied according to evaluation on hourly, daily, and monthly averaged data with RMSE of 161, 80, and 70 , respectively.
While the Heliosat model is a well-known method, there are several other semi-empirical approaches for GHI estimation. [QWL+18] showed that Yang’s hybrid model (YHM) performed best with RMSE of 27.94 when compared to ANN on datasets in China. The physical method YHM relies on the radiation-damping processes, e.g., Rayleigh scattering, aerosol extinction, absorption by ozone, water vaport and permanent gas. These atmospheric variables were derived from MODIS satellite. [QHM+19] presented a mathematical formula of GHI as a function of irradiance with zero surface albedo, global reflectance function of atmosphere and surface albebo, which can be derived from Himawari-8 satellite. The results were validated on 3-year ground station data in Australia with MAE of 55.8 and RMSE of 91.7 . The GHI model with explicit surface albedo was later used in [QHM+21] but all irradiance components were calculated by pre-computing Look Up Table and using radiative transfer model (RTM). The GHI estimation results were validated on data in Australia with an MAE of 54.7 and an RMSE of 90.3 . [BS21] computed the cloud indices using the surface albedo extracted from eight-band imagery of Himawari-8. These cloud indices were converted to clear-sky indices (for all eight bands), and the eight-band clear-sky indices were optimally weighted. The results were validated using data in Singapore, where the scatter plot between the ground-observed clear-sky index and the satellite-estimated values was improved.
@@ 1.2 Site-adaptation techniques
Neural network techniques, presented in the next section, can capture a complex relationship between the satellite data and the corresponding ground-based GHI measurements.
@@ 1.3 Neural network techniques
A methodology that exploits deep-learning or tree-based models to capture a complex relationship between satellite cloud and GHI. To find the best model, we benchmark LSTM, Informer, Transformer, and LightGBM (LGBM) for GHI estimation. Additionally, we improved Informer and Transformer particularly for this application. For LSTM and LGBM, we study the impact of input features. The results show that all these models provide more accurate GHI estimates than the one from a commercial service X. Our implementation are provided at https://github.com/energyCUEE/solarmap-DL-2024.
@@ 2 Thailand solar map
Our Thailand solar map platform is currently running at https://cusolarforecast.com/, where the methodology is a machine learning model that takes input of cloud index, geographical location, and clear-sky irradiance to produce the GHI estimate. The random forest model was implemented from 2022 to June 2024. Our continuing plan of updating the model has been discussed in [SWH+24], and the goal of this paper is to improve the model, considering other tree-based and deep-learning models. The service is currently running under the deep learning model, the Informer, as to be presented subsequently in Section 4.3. This section describes the map requirement and the available input data resources for developing the GHI estimation model.
@@ 3.2 Satellite-derived GHI estimation
When training a model, it is natural to enforce the similarity between the prediction and the ground truth . In this study, we benchmark the following methods: LightGBM (LGBM), LSTM, Informer, Transformer, and the service X. For service X’s predictions, we have downloaded historical GHI estimates at 53 stations during the period of this study. We modified Transformer and Informer to acquire time-series data for solar irradiance forecasting while building upon existing LSTMs with additional linear layers. The hyper-parameters of these models were tuned and set with the values shown in Table 1.
@@ 4.1 Data preparation
The extracted cloud index and ground GHI data are merged into a synchronized 15-min resolution during 03:00-17:00 (local time in Thailand), totaling 1,195,257 samples or 20,924 days. The target forecasting hours range from 7:00 to 17:00. The data is divided with the ratio of 60:20:20 for train:validation:test, where each of the three sets is ensured to contain a uniform mix of three sky conditions see Table 2. To achieve this, we first segmented the data with lags and stacked them as sequences labeled with sky condition. Then, we shuffle the data and split them for training, validation, and testing.
@@ 4.3 Satellite-derived GHI estimation
Figure 11 shows a comparative performance of all GHI estimation models, including LGBM, LSTM, Informer, Transformer, and the service X. All the proposed benchmarked models provide a similar MAEs, that are 20, 78, 98 in clear, partly cloudy, and cloudy sky conditions, respectively, except for the MAE of the service X that is the highest in all conditions; see Figure 11 (Top.). For the MBE, LGBM and the service X have a relatively high negative bias (underestimate) in clear and partly cloudy and with positive bias (overestimate) in the cloudy condition. Meanwhile, LSTM, Informer, and Transformer tend to overestimate—low positive bias in clear and partly cloudy and high positive bias in the cloudy condition, as shown in Figure 11 (Bottom.) The evaluation with the RMSE metric is reported in Table 3. All the proposed models perform almost equally if MAE is considered. However, considering RMSE performance, LGBM has the lowest error in all sky-condition settings. This is because the training loss of the tree model is MSE, while the training loss of deep learning models is set to MAE loss.
| Model | LGBM | LSTM | Informer | Transformer | Service X |
@@
Our GHI estimation at a half-hourly resolution provides the RMSE values (31.54 - 145.74 ) that aligns with previous studies in equatorial regions, i.e., [VPN+23] reported hourly GHI RMSEs of 115-180 across six locations (SAP, TAT, FUA, ISH, DWN, and ASP) using Himawari-8 data; and, [QHM+21] reported an instantaneous GHI RMSE of 118.3 at the DWN station, also using Himawari-8. Our results are in a similar trend to [KBSH18], providing half-hourly GHI RMSEs of 41.7-167 in South Asia (India). [Bri19] reported an average hourly RMSE of 99.5 W/m² in equatorial regions for zenith angles between 0 and 85 degrees (approx. 05:56-18:37), exceeding our target forecasting hours (7:00-17:00). Note that including GHI estimates from early morning and late evening (nearly night-time) in the analysis could potentially underestimate the MAE and RMSE.
@@
The solar irradiance estimation by Informer is provided in Figure 12. The discrepancies between the estimates and ground truth appear to be larger when the sky is cloudy, as shown in Figure 12 (a). Meanwhile, the MAE evaluated for each hour of the day indicates that the model usually suffers during the daytime (10:00-15:00) under the presence of cloud, as shown in Figure 12 (b). The MAE could go up to 140 in this period, and the positive MBE explains that the model mostly overestimates in this period. Figure 12 (c) and (d) show the selected time series with the top three lowest and highest daily average MAE, respectively. The time series with the lowest error often co-occurs with the clear sky condition, where the cloud index is relatively low. The time series with the highest error typically happens with cloud fluctuation during the day. Thus, the model functioned as expected; if CI increases, the estimate decreases. A high estimation error, for example, the middle of the plot (d), could come from approximation in the data processing step. It could be the case that CI was missing before 14:00, and the displayed values were interpolated (as a constant of 0.3 from backward imputation).
@@ 4.4 Solar energy in Thailand
Figure 13(c) illustrates hourly-updated GHI estimates on June 15, 2023. Our GHI estimation model can provide solar energy estimation at the finer temporal resolution, compared to [JMPL13], which was aimed at analyzing a monthly resolution. Our model prioritizes near real-time updates, which offers a significant advantage for operational settings in energy applications where real-time updates are crucial.
@@ 4.5 Deployment and runtime
Our GHI models are deployed to provide the solar map at two spatial resolutions: zoom level 12 (6,843 grids, approximately per grid) and zoom level 14 (93,061 grids, approximately per grid). To efficiently estimate solar irradiance for all grid points simultaneously, we set the batch size equal to the total number of grids in each map. The runtime performance of LSTM, Informer, and Transformer models on an RTX4060 GPU for this process is presented in Table 5. From Section 3.2.1, when directly including latitude and longitude as features for LGBM, the generated solar map displays some abrupt changes in estimated GHI at some region boundary. These unrealistic results can be fixed when removing these two features, with the cost of having 2-3 MAE increased. For this reason, LGBM results are excluded from the runtime report. All models show acceptable performance ( secs. for the largest batch size), with LSTM being the fastest. The Transformer can be an alternative to the Informer because of their similar performance.
@@ 5 Conclusions
This research presented a methodology of a satellite-derived GHI estimation across Thailand with a spatial resolution of , totaling 93,061 grids with 30-minute updates using cloud indices extracted from Himawari-8 satellite imageries, along with the output of Ineichen and Perez’s clear-sky irradiance model as main features of our GHI estimation models. The study explored recently effective machine learning models, including LightGBM (LGBM), LSTM, Informer, and Transformer, to capture complex cloud effects on ground irradiance attenuation. To optimize computational efficiency, input data is processed as time series, with necessary modifications to the Informer and Transformer. Benchmarking against a commercial service from 53 ground stations demonstrates an improved performance, with lower MAE and RMSE values. Our GHI estimates, at a half-hourly resolution and with RMSE values ranging from 31.54 to 142.42 W/m², align with previous studies [KBSH18, Bri19, QHM+21, VPN+23]. We provide real-time solar energy maps at hourly, monthly, and yearly resolutions, and assess computational efficiency at different spatial scales. Our monthly results on March, April, and May are aligned with the monthly average of daily global irradiance in 2013 during the same months [JMPL13]. Additionally, we confirmed the feasibility of using these models for real-time GHI estimation at the required temporal resolution.